OpenStudy (anonymous):

Solve, using the substitution method. d + e = 2 d – e = 4 A. The solution is (3, –1). B. There is no solution. C. The solution is (5, 1). D. There are an infinite number of solutions. I think its A..

OpenStudy (anonymous):

Try plugging in the values-- A. works B. can't be the answer because "A" works C. doesn't work D. There aren't an infinite number of solutions--in general there are only an infinite number of solutions if there are more variables than given equations (i.e. d+e+f = 2, d-e-f=-2), or if one of the functions you have is a multiple of another--i.e something like "d + e = 2, 2d + 2e = 4."

OpenStudy (anonymous):

So A is right..?

OpenStudy (anonymous):

Yeah, A is correct... meant to mention that lol

OpenStudy (anonymous):

okay XD Thank you :)